# Tightening the entropic uncertainty relations with quantum memory in a multipartite scenario

### v1

@@ I Introduction
where $\Delta M=\sqrt{\langle\psi|M^{2}|\psi\rangle-\langle\psi|M|\psi\rangle^{2}}$ is the standard deviation of an observable $M$ and $[M_{1},M_{2}]=M_{1}M_{2}-M_{2}M_{1}$. Here, the lower bound of (1) is state-dependent and becomes trivial when $|\psi\rangle$ is an eigenstate of either $M_{1}$ or $M_{2}$.
@@ II Tripartite QMA-EUR and generalized QMA-EURs for multiple measurements in multipartite systems
We introduce a tripartite QMA-EUR and extend it to the scenario involving $m$ measurements and $n$ memories ($n\leq m$). Consider two arbitrary orthonormal bases $M_{1}=\{|\varphi_{j}^{1}\rangle\}$ and $M_{2}=\{|\varphi_{k}^{2}\rangle\}$ in quantum system $A$ with a finite dimension $d$. Coles et al. Coles and Piani (2014) enhanced the lower bounds of inequalities (3) and (4) as follows,
@@
where $c^{1,2}_{2}$ represents the second largest element of the matrix $[c_{j,k}^{1,2}]$.
@@
with $q(\rho_{A},M_{1},M_{2})=\sum\limits_{j}p^{M_{1}}_{j}\log_{2}(1/\max\limits_{k}c^{1,2}_{jk})$ and $q(\rho_{A},M_{2},M_{1})=\sum\limits_{k}p^{M_{2}}_{k}\log_{2}(1/\max\limits_{j}c^{1,2}_{jk})$, where $p^{M_{1}}_{j}=\langle\varphi_{j}^{1}|\rho_{A}|\varphi_{j}^{1}\rangle$ and $p^{M_{2}}_{k}=\langle\varphi_{k}^{2}|\rho_{A}|\varphi_{k}^{2}\rangle$. $q^{M_{1,2}}$ is defined as the minimum value taken over all possible states $\rho_{A}$ Coles and Piani (2014),
@@
where $\lambda_{\min}[\cdot]$ denotes the smallest eigenvalue, and $\Delta(p)=p\Delta_{M_{1}M_{2}}+(1-p)\Delta_{M_{2}M_{1}}$. Here, $\Delta_{M_{1}M_{2}}=\sum\limits_{j}\log_{2}(1/\max\limits_{k}c^{1,2}_{jk})|\varphi_{j}^{1}\rangle\langle\varphi_{j}^{1}|$ and $\Delta_{M_{2}M_{1}}=\sum\limits_{k}\log_{2}(1/\max\limits_{j}c^{1,2}_{jk})|\varphi_{k}^{2}\rangle\langle\varphi_{k}^{2}|$. Note that $Q^{M_{1,2}}\geq q^{M_{1,2}}_{MU}$ and the inequality becomes an equality when $d=2$ Coles and Piani (2014).
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{-1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)+\sum_{t}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A|B_{t})+\max\{0,\delta_{mn}\},$ (12)
@@
$\delta_{mn}=$ $\frac{m(m-1)-\sum_{t=1}\limits^{n}m_{t}(m_{t}-1)}{2(m-1)}S(A)+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})-\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t})$
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{-1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)+\sum_{t}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A|B_{t})+\max\{0,\delta_{mn}^{\prime}\}$ (13)
@@
$\delta_{mn}^{\prime}=$ $\frac{1}{m-1}\log_{2}\frac{\left(\prod\limits_{i<j}^{m}c_{i,j}\right)}{b^{m-1}}+(m-1)S(A)-\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t})$
$-\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A)+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t}),$
@@
Following the approach in Zhang and Fei (2023); Wehner and Winter (2010); Schwonnek (2018), we derive a simply constructed bound (SCB) and a tighter lower bound for QMA-EUR concerning $m$ measurements within the framework of $n$ memories.
Theorem 2.
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})+\max\{0,\delta_{mn}\},$ (14)
@@
For all $i<j$, we obtain $\frac{m(m-1)}{2}$ similar inequalities. Summarizing these inequalities and dividing both sides by $m-1$, we get
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{m}{2}S(A)-\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t}).$ (15)
@@
By summing over these inequalities for all $i<j$ and subsequently dividing both sides of the summed inequality by $m-1$, we obtain
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t}),$ (16)
@@
Theorem 3.
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\sum_{t}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})+\max\{0,\delta_{mn}^{{}^{\prime\prime}}\},$ (17)
@@
$\delta_{mn}^{{}^{\prime\prime}}=$ $-\frac{1}{m}\omega\beta+(m-1)S(A)-\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A)+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})$
@@
By employing majorization theory and actions of the symmetric group, Xiao et al. Xiao et al. (2016) presented an admixture bound of the entropic uncertainty relation for multiple measurements:
$\sum_{i=1}^{m}H(M_{i})\geq-\frac{1}{m}\omega\beta+(m-1)S(A),$ (19)
where $\omega$ denotes the universal majorization bound for $m$ measurements, and $\beta$ is a vector of logarithmic distributions. It is notable that this admixture bound surpasses the result achieved by Liu et al. Liu et al. (2015). Integrating the inequality (19) with equation (18) we derive a more stringent lower bound,
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $-\frac{1}{m}\omega\beta+(m-1)S(A)-\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t}).$ (20)
@@
We now compare our Theorems 2 and 3 with the results given by Zhang et al. Zhang and Fei (2023). For convenience, we denote the right-hand sides of the inequalities (12), (13), (14) and (17) by $Lb1$, $Lb2$, $LB1$ and $LB2$, respectively. Through straightforward calculation, we obtain that
$LB1-Lb1=\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)\geq 0,$
in which the equality holds when the state in system $A$ is a qubit Coles and Piani (2014). Moreover,
$LB2-Lb2=$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)+\max\{0,\delta_{mn}^{{}^{\prime\prime}}\}-\max\{0,\delta_{mn}^{\prime}\}$
$=$ $-\frac{1}{m}\omega\beta-(-\log_{2}b)>0,$
@@
where $b=\max\limits_{k_{m}}\{\sum\limits_{k_{2}\sim k_{m-1}}\max\limits_{k_{1}}|\langle\varphi_{k_{1}}^{1}|\varphi_{k_{2}}^{2}\rangle|^{2}\prod_{i=2}^{m-1}|\langle\varphi_{k_{i}}^{i}|\varphi_{k_{i+1}}^{i+1}\rangle|^{2}\}$ is the same as the one given in inequality (13), and the second equation holds when $\delta_{mn}^{{}^{\prime\prime}},\delta_{mn}^{\prime}\geq 0$. In Xiao et al. (2016) Xiao et al. presented an admixture lower bound that surpassed the result achieved by Liu et al.’s Liu et al. (2015), which ensures that $LB2-lb2$ is positive.
@@
To further illustrate the effectiveness of our results, we consider three measurements $M_{1}$, $M_{2}$ and $M_{3}$ in a two-dimensional Hilbert space with eigenvectors $|\psi^{1}_{1}\rangle=\displaystyle\frac{1}{\sqrt{2}}(1,1)^{T}$, $|\psi^{1}_{2}\rangle=\displaystyle\frac{1}{\sqrt{2}}(1,-1)^{T}$, $|\psi^{2}_{1}\rangle=\displaystyle\frac{1}{2}(1,\sqrt{3})^{T}$, $|\psi^{2}_{2}\rangle=\displaystyle\frac{1}{2}(\sqrt{3},-1)^{T}$, $|\psi^{3}_{1}\rangle=(0,1)^{T}$ and $|\psi^{3}_{2}\rangle=(1,0)^{T}$. By computation we obtain $-\frac{1}{m}\omega\beta=0.827$ and $-\log_{2}b=0.5$. That is to say, $LB2-lb2=0.327>0.$ A more general example involves three measurements, $M_{1}$, $M_{1}$ and $M_{3}$ in a three-dimensional Hilbert space with the following eigenvectors $|\varphi^{1}_{1}\rangle=(1,0,0)^{T}$, $|\varphi^{1}_{2}\rangle=(0,1,0)^{T}$, $|\varphi^{1}_{3}\rangle=(0,0,1)^{T}$, $|\varphi^{2}_{1}\rangle=\displaystyle(\frac{1}{\sqrt{2}},0,-\frac{1}{\sqrt{2}})^{T}$, $|\varphi^{2}_{2}\rangle=(0,1,0)^{T}$, $|\varphi^{2}_{3}\rangle=\displaystyle(\frac{1}{\sqrt{2}},0,\frac{1}{\sqrt{2}})^{T}$, $|\varphi^{3}_{1}\rangle=(\sqrt{a},e^{i\phi}\sqrt{1-a},0)^{T}$, $|\varphi^{3}_{2}\rangle=(\sqrt{1-a},-e^{i\phi}\sqrt{a},0)^{T}$ and $|\varphi^{3}_{3}\rangle=(0,0,1)^{T}$. As illustrated in Fig.1, our bound consistently outperforms previous bounds for $\phi=\frac{\pi}{2}$ and $10<a<$ 1.

### v2

@@ Abstract
Keywords: Quantum memory; Entropic uncertainty relations; Unilateral coherence; Quantum key distribution
@@ I Introduction
where $\Delta M=\sqrt{\langle\psi|M^{2}|\psi\rangle-\langle\psi|M|\psi\rangle^{2}}$ is the standard deviation of an observable $M$ and $M_{1},M_{2}]=M_{1}M_{2}-M_{2}M_{1}$. Here, the lower bound of (1) is state-dependent and becomes trivial when $|\psi\rangle$ is an eigenvector of either $M_{1}$ or $M_{2}$.
@@ II Tripartite QMA-EUR and generalized QMA-EURs for multiple measurements in multipartite systems
We introduce a tripartite QMA-EUR and extend it to the scenario involving $m$ measurements and $n$ memories ($n\leq m$). Consider two measurements $M_{1}$ and $M_{2}$ given by eigenvectors $\{|\varphi_{j}^{1}\rangle\}$ and $\{|\varphi_{k}^{2}\rangle\}$, respectively, in a quantum system $A$ with finite dimension $d$. Coles et al. Coles and Piani (2014) enhanced the lower bounds of inequalities (3) and (4) as follows,
@@
where $c^{1,2}_{2}$ represents the second largest element of the matrix $c_{j,k}^{1,2}]$.
@@
with $q(\rho_{A},M_{1},M_{2})=\sum\limits_{j}p^{M_{1}}_{j}\log_{2}(1/\max\limits_{k}c^{1,2}_{jk})$ and $q(\rho_{A},M_{2},M_{1})=\sum\limits_{k}p^{M_{2}}_{k}\log_{2}(1/\max\limits_{j}c^{1,2}_{jk})$, where $p^{M_{1}}_{j}=\langle\varphi_{j}^{1}|\rho_{A}|\varphi_{j}^{1}\rangle$, $p^{M_{2}}_{k}=\langle\varphi_{k}^{2}|\rho_{A}|\varphi_{k}^{2}\rangle$. $q^{M_{1,2}}$ is defined as the minimum value taken over all possible states $\rho_{A}$ Coles and Piani (2014),
@@
where $\lambda_{\min}[\cdot]$ denotes the smallest eigenvalue, and $\Delta(p)=p\Delta_{M_{1}M_{2}}+(1-p)\Delta_{M_{2}M_{1}}$. Here, $\Delta_{M_{1}M_{2}}=\sum\limits_{j}\log_{2}(1/\max\limits_{k}c^{1,2}_{jk})|\varphi_{j}^{1}\rangle\langle\varphi_{j}^{1}|$ and $\Delta_{M_{2}M_{1}}=\sum\limits_{k}\log_{2}(1/\max\limits_{j}c^{1,2}_{jk})|\varphi_{k}^{2}\rangle\langle\varphi_{k}^{2}|$. Note that $Q^{M_{1,2}}\geq q^{M_{1,2}}_{MU}=\log_{2}(1/c_{\max}^{1,2})$ for almost all choices of the eigenvectors of $M_{i}$ and $M_{j}$. For $Q^{M_{1,2}}$ is one of $\widetilde{q}^{M_{1,2}}$ and $q^{M_{1,2}}(\rho_{A})$ the inequality becomes an equality when the system $A$ is a qubit ($d=2$), and if we take $Q^{M_{1,2}}=q^{M_{1,2}}$, the equation holds when the function $\lambda_{\min}[\Delta(p)]$ is independent of $p$ Coles and Piani (2014).
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $-\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)$ (12)
$+\sum_{t}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A|B_{t})$ $+\max\{0,\delta_{mn}\},$
@@
$\delta_{mn}=$ $\frac{m(m-1)-\sum_{t=1}\limits^{n}m_{t}(m_{t}-1)}{2(m-1)}S(A)$
$+$ $\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})$
- $\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t})$
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $-\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)$ (13)
$+\sum_{t}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A|B_{t})$ $+\max\{0,\delta_{mn}^{\prime}\}$
@@
$\delta_{mn}^{\prime}=$ $\frac{1}{m-1}\log_{2}\frac{\left(\prod\limits_{i<j}^{m}c_{i,j}\right)}{b^{m-1}}+(m-1)S(A)$
- $\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A)+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})$
- $\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t}),$
@@
Following the approach in Zhang and Fei (2023); Wehner and Winter (2010); Schwonnek (2018), we derive two tighter lower bounds for QMA-EURs concerning $m$ measurements within the framework of $n$ memories.
Theorem 2
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}$ (14)
$+$ $\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})$
$+$ $\max\{0,\delta_{mn}\},$
@@
For all $i<j$, we obtain $\frac{m(m-1)}{2}$ similar inequalities. Summarizing these inequalities and dividing both sides by m1m-1, we get
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{m}{2}S(A)$ (15)
- $\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t}).$
@@
By summing these inequalities for all $i<j$ and subsequently dividing both sides of the summed inequality by m1m-1, we obtain
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}$ (16)
$+$ $\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t}),$
@@
Let $\rho$ be a mixed state in a $d$ dimensional Hilbert space, and $M_{i}$ $(i=1,2,\ldots,m)$ be $m$ measurements with a set of orthonormal eigenvectors $\{|\varphi^{i}_{i_{m}}\rangle\}$ $(i_{m}=1,2,\ldots,d)$. Denote $p^{i}_{i_{m}}=\langle\varphi^{i}_{i_{m}}|\rho|\varphi^{i}_{i_{m}}\rangle$ the probability distributions obtained by measuring $\rho$ with respect to bases $\{|\varphi^{i}_{i_{m}}\rangle\}$. By employing majorization theory and actions of the symmetric group, Xiao et al. Xiao et al. (2016) presented an admixture bound of the entropic uncertainty relation for multiple measurements:
$\sum_{i=1}^{m}H(M_{i})\geq-\frac{1}{m}\omega\beta+(m-1)S(A),$ (17)
where $\beta$ is the $d^{m}$ dimension vector $(\log(\omega\cdot\mathfrak{U}_{i_{1},i_{2}\cdots i_{m}}))^{\downarrow}$ defined as follows. The symbol $\downarrow$ stands for arranging the components in descending order. For each multi-index $(i_{1},i_{2}\cdots,i_{m})$, the $d^{m}$ dimensional vector $\mathfrak{U}_{i_{1},i_{2}\cdots i_{m}}$ has entries of the form $c(1,2,\cdots m)c(2,3,\cdots,1)\cdots c(m,1,\cdots,m-1)$ sorted in decreasing order with respect to the indices $(i_{1},i_{2},\cdots,i_{m})$, where $c(1,2,\cdots m)=\sum_{i_{2},\cdots,i_{m-1}}\max\limits_{i_{1}}c(\varphi_{i_{1}}^{1},\varphi_{i_{2}}^{2})\cdots c(\varphi_{i_{m-1}}^{m-1},\varphi_{i_{m}}^{m})$. The majorization vector bound $\omega$ for probability tensor distributions $(p^{1}_{i_{1}}p^{2}_{i_{2}}\cdots p^{m}_{i_{m}})_{i_{1},i_{2},\cdots,i_{m}}$ of a state $\rho$ is the $d^{m}$-dimensional vector $\omega=(\Omega_{1},\Omega_{2}-\Omega_{1},\cdots,1-\Omega_{a},0,\cdots,0)$, where
$\Omega_{k}=(\frac{s_{k}}{m})^{m}.$
We have $\Omega_{1}\leq\Omega_{2}\cdots\Omega_{a}<1$ for some integer $a\leq d^{m}-1$ with $\Omega_{a+1}=1$. The more detailed definition of $s_{k}$, see Xiao et al. (2016).
Theorem 3
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}$ (18)
$+$ $\frac{1}{m-1}\sum_{t}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})$
$+$ $\max\{0,\delta_{mn}^{{}^{\prime\prime}}\},$
@@
$\delta_{mn}^{{}^{\prime\prime}}=$ $-\frac{1}{m}\omega\beta+(m-1)S(A)-\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A)$
$+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})$
@@
Integrating the inequality (17) with equation (19) we derive a more stringent lower bound,
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $-\frac{1}{m}\omega\beta+(m-1)S(A)$ (20)
$-\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}\mathcal{I}(M_{i}:B_{t}).$
@@
It is interesting to compare our Theorems 2 and 3 with the results given by Zhang et al. Zhang and Fei (2023) and compare the lower bounds of Theorems 2 and 3. For convenience, we denote the right-hand sides of the inequalities (12), (13), (14) and (18) by $lb1$, $lb2$, $LB1$ and $LB2$, respectively. Through straightforward calculation, we obtain that
$LB1-LB2=$ $\max\{0,\delta_{mn}\}-\max\{0,\delta_{mn}^{{}^{\prime\prime}}\}$
$=$ $\frac{2-m}{2}S(A)+\frac{1}{m}\omega\beta+\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}.$
$LB1-lb1=$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right),$
$LB1-lb2=$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)+$
$\max\{0,\delta_{mn}\}-\max\{0,\delta_{mn}^{\prime}\}$
$=$ $\frac{2-m}{2}S(A)+\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\log_{2}b,$
$LB2-lb1=$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)$
$+$ $\max\{0,\delta_{mn}^{{}^{\prime\prime}}\}-\max\{0,\delta_{mn}\}$
$=$ $\frac{m-2}{2}S(A)+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)-\frac{1}{m}\omega\beta$
and
$LB2-lb2=$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{1}{m-1}\log_{2}\left(\prod_{i<j}^{m}c_{i,j}\right)$
$+$ $\max\{0,\delta_{mn}^{{}^{\prime\prime}}\}-\max\{0,\delta_{mn}^{\prime}\}$
$=$ $-\frac{1}{m}\omega\beta+\log_{2}b,$
where
$b=\max\limits_{k_{m}}\{\sum\limits_{k_{2}\sim k_{m-1}}\max\limits_{k_{1}}|\langle\varphi_{k_{1}}^{1}|\varphi_{k_{2}}^{2}\rangle|^{2}\prod_{i=2}^{m-1}|\langle\varphi_{k_{i}}^{i}|\varphi_{k_{i+1}}^{i+1}\rangle|^{2}\}$
@@
is the same as in inequality (13), and the last equalities in the above equations hold when $\delta_{mn},\delta_{mn}^{{}^{\prime\prime}},\delta_{mn}^{\prime}\geq 0$. In Xiao et al. (2016), Xiao et al. presented an admixture lower bound that surpassed the result achieved by Liu et al.’s Liu et al. (2015), which ensures that LB2lb2LB2-lb2 is non-negative. Combining with $Q^{M_{i,j}}\geq q^{M_{i,j}}_{MU}$, we conclude that our lower bounds $LB1$ and $LB2$ are tighter than $lb1$ and $lb2$, respectively.
Two groups of orthonormal bases $\{|\varphi^{i}_{i_{m}}\rangle\}$ and $\{|\varphi^{j}_{i_{n}}\rangle\}$ in a $d$-dimension Hilbert space are called mutually unbiased if $|\langle\varphi^{i}_{i_{m}}|\varphi^{j}_{i_{n}}\rangle|^{2}=1/d$ holds for all $i_{m}$ and $i_{n}$Durt et al. (2010); Designolle et al. (2019). In fact, when we consider mutually unbiased measurements (MUMs), we obtain $Q^{M_{i,j}}=q^{M_{i,j}}_{MU}$ and $\omega\beta=-m\log_{2}d$, i.e., $LB1=lb1$, $LB2=lb2$. We also find that
$lb1-lb2=$ $\delta_{mn}^{\prime}-\delta_{mn}$
$=$ $\frac{m-2}{2}(\log_{2}d-S(A))\geq 0,$
$LB1-LB2=$ $\frac{2-m}{2}S(A)+\frac{1}{m}\omega\beta+\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}$
$=$ $\frac{2-m}{2}S(A)+\frac{1}{m}\omega\beta+\frac{m}{2}\log_{2}d$
$=$ $\frac{m-2}{2}(\log_{2}d-S(A))\geq 0.$
The above inequalities are saturated when $m=2$ or $\log_{2}d=S(A)$ and are strictly larger for $m\geq 3$.
To further illustrate the effectiveness of our results, we present the following examples and consider the values of LB1lb1LB1-lb1, LB1lb2LB1-lb2, LB2lb1LB2-lb1, LB2lb2LB2-lb2 and LB1LB2LB1-LB2. For convenience, we take $Q^{M_{i,j}}=\widetilde{q}^{M_{1,2}}$ in the examples of this paper.
@@
Example 1 Consider the three-qutrit state $|\Phi\rangle=\displaystyle\frac{1}{\sqrt{3}}(|000\rangle+|111\rangle+|222\rangle)$ and three measurements $M_{1}$, $M_{2}$ and $M_{3}$ in a three-dimensional Hilbert space with the following eigenvectors, $|\varphi^{1}_{1}\rangle=(\sqrt{a},\sqrt{1-a},0)^{T}$, $|\varphi^{1}_{2}\rangle=(e^{\mathrm{i}\phi}\sqrt{1-a},-e^{\mathrm{i}\phi}\sqrt{a},0)^{T}$, $|\varphi^{1}_{3}\rangle=(0,0,1)^{T}$, $|\varphi^{2}_{1}\rangle=\frac{1}{\sqrt{3}}(1,1,1)^{T}$, $|\varphi^{2}_{2}\rangle=\frac{1}{\sqrt{3}}(1,\omega^{\ast},\omega)^{T}$, $|\varphi^{2}_{3}\rangle=\frac{1}{\sqrt{3}}(1,\omega,\omega^{\ast})^{T}$, $|\varphi^{3}_{1}\rangle=\frac{1}{\sqrt{3}}(1,1,\omega^{\ast})^{T}$, $|\varphi^{3}_{2}\rangle=\frac{1}{\sqrt{3}}(1,\omega,\omega)^{T}$, $|\varphi^{3}_{3}\rangle=\frac{1}{\sqrt{3}}(1,\omega^{\ast},1)^{T}$, where $\omega=e^{\frac{2}{3}\pi\mathrm{i}}$, $\omega^{\ast}$ represents the conjugate of $\omega$. For the convenience of calculation, we take $a=0.95$ , $\phi=\pi$ and obtain LB1LB2=0.0736LB1-LB2=0.0736, LB1lb1=0.0366LB1-lb1=0.0366, LB1lb2=0.4366LB1-lb2=0.4366, LB2lb1=0.0370LB2-lb1=-0.0370, LB2lb2=0.3630LB2-lb2=0.3630. It can be seen that for this example, $LB1>lb1>LB2>lb2$, i.e., our lower bound $LB1$ is tighter than the previous results $lb1$ and $lb2$ and $LB1>LB2$.
Example 2 Consider the three-qutrit $|\Phi\rangle=\displaystyle\frac{1}{\sqrt{3}}(|012\rangle+|120\rangle+|201\rangle)$ and three measurements $M_{1}$, $M_{1}$ and $M_{3}$ in a three-dimensional Hilbert space with the following eigenvectors, $|\varphi^{1}_{1}\rangle=(1,0,0)^{T}$, $|\varphi^{1}_{2}\rangle=(0,0,1)^{T}$, $|\varphi^{1}_{3}\rangle=(0,-1,0)^{T}$, $|\varphi^{2}_{1}\rangle=(-\frac{1}{3},-\frac{2}{3},\frac{2}{3})^{T}$, $|\varphi^{2}_{2}\rangle=(-\frac{2}{\sqrt{5}},\frac{1}{\sqrt{5}},0)^{T}$, $|\varphi^{2}_{3}\rangle=(\frac{2}{3\sqrt{5}},\frac{4}{3\sqrt{5}},\frac{5}{3\sqrt{5}})^{T}$, $|\varphi^{3}_{1}\rangle=(\sqrt{a},\sqrt{1-a},0)^{T}$, $|\varphi^{3}_{2}\rangle=(e^{\mathrm{i}\phi}\sqrt{1-a},-e^{\mathrm{i}\phi}\sqrt{a},0)^{T}$, $|\varphi^{3}_{3}\rangle=(0,0,1)^{T}$. For the convenience of calculation, we take $\phi=\pi$, $1\geq a\geq 0$ and obtain the following Figures.
Figure 1: The horizontal axis represents the parameter $a$, while the vertical axis represents the difference (diff) between our results and the previous results. The green (circle), black (solid), blue (triangle) and the red (square) curves represent the values of LB2lb1LB2-lb1, LB2lb2LB2-lb2, LB1lb2LB1-lb2 and LB1LB2LB1-LB2, respectively.